OpenStudy (anonymous):

y^2+6y+8x+25=0 find the vertex focus and directrix of the parabola

OpenStudy (anonymous):

first completing squares

OpenStudy (anonymous):

y^2 + 6y + 9 = 9 -25 - 8x (y + 3)^2 = -16 - 8x (y + 3)^2 = -8 ( 2 + x)

OpenStudy (anonymous):

from here we can see that vertex is ( -2 ; -3)

OpenStudy (anonymous):

would P = -2?

OpenStudy (anonymous):

4p = -8 yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

it was a pleasure to help you

OpenStudy (anonymous):

So what conic section would this be?

OpenStudy (anonymous):

parabola